# Illusion or Algorithm? Investigating Memorization, Emergence, and Symbolic Processing in In-Context Learning

### v2

@@ Abstract
Large-scale Transformer language models (LMs) trained solely on next-token prediction with web-scale data can solve a wide range of tasks after seeing just a few examples. The mechanism behind this capability, known as in-context learning (ICL), remains both controversial and poorly understood. Some studies argue that it is merely the result of memorizing vast amounts of data, while others contend that it reflects a fundamental, symbolic algorithmic development in LMs. In this work, we introduce a suite of investigative tasks and a novel method to systematically investigate ICL by leveraging the full Pythia scaling suite, including interim checkpoints that capture progressively larger amount of training data. By carefully exploring ICL performance on downstream tasks and simultaneously conducting a mechanistic analysis of the residual stream’s subspace, we demonstrate that ICL extends beyond mere “memorization” of the training corpus, yet does not amount to the implementation of an independent symbolic algorithm. Our results also clarify several aspects of ICL, including the influence of training dynamics, model capabilities, and elements of mechanistic interpretability. Overall, our work advances the understanding of ICL and its implications, offering model developers insights into potential improvements and providing AI security practitioners with a basis for more informed guidelines. [^1] The code and data of this work are publicly available online: https://github.com/UKPLab/arxiv2025-icl-investigation.
@@ 1 Introduction
Large-scale Transformer based language models (LMs) trained exclusively on the next-token prediction task acquire the ability to perform a novel task based solely on a few demonstrations provided within the prompt, without any gradient updates—a phenomenon known as in-context learning (ICL; Brown et al., 2020). ICL, also referred to as few-shot learning, is a fundamental mechanism in Language Models (LMs) (Lampinen et al., 2024). Therefore, understanding the nature and generalization limits of ICL during the the next-token prediction pre-training process, is crucial. Additionally, understanding the nature of ICL is central to the safe and secure deployment of generative AI, as it has been argued that since increased scale alone is known to unlock new abilities in LMs, further scaling up has the potential to unlock latent hazardous abilities which carry important security and safety implications. Simultaneously, Lu et al. (2024) provide strong correlations between the ability of a model to solve tasks in few-shot ICL and their generalizability on novel tasks upon instruction tuning. Therefore, understanding the fundamental mechanisms of ICL has broad implications, including safety, security, and the capabilities of user-facing, instruction-tuned LMs.
Despite its importance, the current understanding of ICL in LMs remains limited and is often clouded by contradictory claims and findings, unlike our understanding of traditional in-weight learning via gradient descent. Specifically, researchers continue to debate whether ICL is functionally equivalent to gradient descent (Von Oswald et al. (2023) vs. Deutch et al. (2024)), and whether ICL is primarily regurgitation data and patterns encountered during pre-trained or genuine generalization—even to the point of implementing symbolic algorithms (Golchin et al. (2024) vs. Elhage et al. (2021)), inter alia. A significant majority of prior attempts to investigate the characteristics of ICL focus on either Transformers trained on purely synthetic data (Edelman et al., 2024; Singh et al., 2023) or mechanism construction within toy-sized Transformers (Elhage et al., 2021; Olsson et al., 2022); as such, it is hard to scale these findings to LMs trained on real-world language data. Simultaneously, there has been a discrepancy in the tasks considered to characterize ICL; e.g., Golchin et al. (2024) consider different NLP benchmark tasks (with the possibility of leakage), while Olsson et al. (2022) define ICL more narrowly as an LM’s ability to replicate patterns from context. Intuitively, while the latter is a prerequisite for the former, several missing intermediary elements prevent a direct equivalence, leading to contradictory outcomes.
@@
We find that when provided with in-context examples, LMs trained on real-world corpora display a peculiar mix of generalized pattern-matching abilities on the one hand, and a dependence on pattern statistics on the other. Specifically, they can detect and follow patterns presented in-context using random token orderings—and therefore are unlikely to have memorized them during pre-training—refuting prior claims by Golchin et al.’s (2024). However, this pattern-learning ability declines when statistically rarer tokens are used to construct the patterns or when the pattern statistics in the examples are obfuscated, directly contradicting prior claims by Olsson et al.’s (2022), who described ICL as symbolic algorithm execution. These two findings significantly refine our understanding of the mechanistic basis for ICL in LM, with implications to interpretability and circuit discovery (Wang et al., 2022; Conmy et al., 2023; Yu et al., 2024) in LMs, as contemporary approaches often operate under the assumption that symbolic functions are being implemented within the models’ internals.
@@
Upon investigating the model internals, we identify an aspect of the development that remains stable across different ICL characteristics. Specifically, we show that the dimensionality of the subspace of the residual stream (Elhage et al., 2021) used by the model to “carry” the answer token follows a consistent pattern: an initial rise in subspace allocation, corresponding to a “general” development of ICL competence, followed by a decay that saturates over training. Moreover, development of ICL competence through pre-training correlates with the model learning to use a common subspace across tasks. Both of these developments follow a predictable pattern, serving as a predictable reference variable for ‘occasionally abrupt’ development of ICL competence.
All of these findings bear importance to the notion of “emergent abilities of Large Language Models” introduced by Wei et al. (2022) and Ganguli et al. (2022) which are also related to ICL as described by Olsson et al. (2022); Lu et al. (2024): the predictability, or the lack thereof, of ICL has direct relevance to the possibility of other abilities being unexpectedly unlocked with further increase in scale, with significant implications to AI security.
@@ 2 Related Work
Brown et al. (2020) first empirically demonstrated that LMs, trained purely on next-token prediction (a.k.a. the language modelling task), can learn to perform a task by observing a few demonstrations in the input prompt. In this work, we coin the term ICL competence to refer to the general prowess of the LM in performing ICL, quantified by the model’s performance on various ICL tasks. Our notion of ICL competence is much similar to Lampinen et al. (2024) who argue that ICL encapsulate a broad range of tasks beyond supervised few-shot learning. The number of studies showcasing what tasks LMs can do from ICL examples is too many to discuss individually within the scope of this paper; instead, we refer to the elaborate survey by Dong et al. (2024a). Min et al. (2022) attempted to disentangle the factors influencing an LM’s ICL competence by perturbing the input-output space, subsequently claiming that LMs are often able to perform in-context classification correctly even when the labels are randomly permuted in the context. Wei et al. (2023) showcased that semantic priors of input-label mapping play a vital role in smaller LMs, while large LMs can perform ICL with semantically unrelated labels. Shi et al. (2024) explained this effect of scale on ICL competence by larger LMs’ ability to process more in-context features. These works remain primarily focused on the static abilities of the trained model, without investigating how these abilities might have arisen via large-scale pre-training. In a somewhat orthogonal direction, Akyürek et al. (2024) investigated ICL competence through the lens of formal language modelling.
@@
A distinct line of attempts to understand ICL presents constructions of Transformer attention that can simulate gradient descent to argue that Transformers can implement certain classes of learning algorithms implicitly as ICL (Akyürek et al., 2022; Von Oswald et al., 2023). (Deutch et al., 2024) argued against this expressiveness-based line of reasoning and pointed towards the distributional properties of training data that possibly dictate the development of ICL. This debate has been formalized to the question of whether ICL and gradient descent are equivalent or not (Dai et al., 2023; Mahdavi et al., 2024; Deutch et al., 2024). Prior work by Chan et al. (2022) has also pointed towards the data-distributional sources of ICL. Singh et al. (2023) demonstrated the transient nature of ICL in small Transformer models trained on synthetic data. (Chan et al., 2025) confirmed this transient nature and provided theoretical results that explain the learning dynamics of ICL. These works primarily identify the long-tailed distribution of the data as being the driver of ICL, whereas Bratuli’c et al. (2025) claimed that conceptual repetitions in the pre-training data are the primary determiner of ICL. Most of these prior investigations primarily rely on toy-sized Transformers and/or synthetic pre-training data. On the other hand, we focus on Transformers of a wide range of parameter scales, trained on real web-scale natural language data. To the best of our knowledge, Wang et al. (2024) comes as the only similar prior work in this direction; however, they sought to identify the dynamics of task recognition from ICL examples vs. task learning as they progress through pre-training.
@@
Elhage et al. (2021) and Olsson et al. (2022) were among the first to demonstrate that LMs can consistently complete previously seen patterns in-context: [A][B]...[A]→[B]. They conjectured that certain internal model mechanisms (specifically, induction heads) form to perform this pattern-completion task as a symbolic algorithm. This hypothesis has since underpinned mechanistic interpretability research, which assumes LMs develop internal mechanisms akin to symbolic algorithms and tasks researchers with identifying them. It has inspired projects like Tracr (Lindner et al., 2023), which compiles symbolic algorithms into transformer weights, and circuit discovery methods (Wang et al., 2022; Conmy et al., 2023) to uncover them. Cho et al. (2025) recently identified the mechanistic circuitry of ICL as a three-step process. Upon analysis on synthetically trained Transformers, Park et al. (2024) found that the development of ICL is a juxtaposition of multiple algorithms and their completion, and any claim of ‘singular’ ICL mechanism should be taken with caution. In a somewhat less fine-grained method of interpreting ICL, the notion of function vectors (Todd et al., 2024) or task vectors (Todd et al., 2023) has been proposed.  Yin & Steinhardt (2025) identified attention heads that construct function vectors and show that 1) these heads are formed from induction heads upon further pre-training, and 2) these heads are the fundamental driver of ICl-specific circuits and not induction heads, as presumed by Elhage et al. (2021). In line with the school of mechanistic interpretability, our work validates the algorithmic nature of ICL, but puts significant limitations on the notion as well. While we seek to identify signals internal to the model that implicate ICL and its development with training and parameter scaling (Section 6), we do not focus on recovering the exact neural algorithmic implementation of ICL.
@@
Table 1: An overview of our six proposed tasks: lsc, lscg, wc, wi, tt, and cf. In the examples, the prompt text is highlighted in green, and the expected target token is highlighted in orange.
@@
Literal Sequence Copying (LSC): We start with providing a systematic empirical evaluation of the literal sequence copying phenomenon identified by Olsson et al. (2022). In particular, we aim to test whether the model can generate T when prompted with P T R P, where T is a target token and P and R are sequences of randomly sampled tokens of length ||||P|||| and ||||R||||, respectively.
@@
Task Configurations: Pattern length (||||P||||): the number of tokens in the literal sequence pattern prefix (P). Random gap length (||||R||||): the number of tokens in the random gap sequence (R).
@@
Template: P U X T R P V X T Example: $\overbrace{{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{Category 40 ids \text@underline{Garlic right} node struction}}}}}^{\text{Pattern w/ RandomGap: }{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{P}}}}^{*}{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{U}}}}^{*}{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{X}}}}{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{T}}}}}{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{ }}}}\overbrace{{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{Yolk yes Production}}}}}^{\text{Random Gap: }{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{R}}}}^{*}}$ $\underbrace{{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{Category 40 ids \text@underline{total Content} node}}}}}_{\text{Pattern w/ a Different Random Gap: }{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{P}}}}^{*}{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{V}}}}^{*}{\color[rgb]{0,0.3671875,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3671875,0.09765625}\textbf{\scriptsize{{X}}}}}{\color[rgb]{1,0.30078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.30078125,0}\textbf{\scriptsize{{ }}}}\underbrace{{\color[rgb]{1,0.30078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.30078125,0}\textbf{\scriptsize{{struction}}}}}_{\text{Target: }{\color[rgb]{1,0.30078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.30078125,0}\textbf{\scriptsize{{T}}}}}$
Task Configurations: Pattern length (||||P||||): the number of tokens in the literal sequence pattern prefix (P). Random gap length (||||R||||): the number of tokens in the random gap sequence (R). Random gap in pattern length (||||G||||): the number of tokens in the random gap sequence in the pattern with random gap (U or U).
@@
Task Configurations: Number of features (||||F||||): the number of tokens to identify in sequence. Number of labels (||||L||||): the number classification labels. Number of distractors (||||D||||): the number of distractor tokens in the classification sequence.
@@
Example: If we swtich the capital of Canada and Germany, then Canada’s capital is Berlin and Germany’s capital is Ottawa

### v3

@@ Abstract
Large-scale Transformer language models (LMs) trained solely on next-token prediction with web-scale data can solve a wide range of tasks after seeing just a few examples. The mechanism behind this capability, known as in-context learning (ICL), remains both controversial and poorly understood. Some studies argue that it is merely the result of memorizing vast amounts of data, while others contend that it reflects a fundamental, symbolic algorithmic development in LMs. In this work, we introduce a suite of investigative tasks and a novel method to systematically investigate ICL by leveraging the full Pythia scaling suite, including interim checkpoints that capture progressively larger amount of training data. By carefully exploring ICL performance on downstream tasks and simultaneously conducting a mechanistic analysis of the residual stream’s subspace, we demonstrate that ICL extends beyond mere “memorization” of the training corpus, yet does not amount to the implementation of an independent symbolic algorithm. Our results also clarify several aspects of ICL, including the influence of training dynamics, model capabilities, and elements of mechanistic interpretability. Overall, our work advances the understanding of ICL and its implications, offering model developers insights into potential improvements and providing AI security practitioners with a basis for more informed guidelines. [^1] The code and data of this work are publicly available online: https://github.com/UKPLab/tmlr2025-icl-investigation.
@@ 1 Introduction
Large-scale Transformer based language models (LMs) trained exclusively on the next-token prediction task acquire the ability to perform a novel task based solely on a few demonstrations provided within the prompt, without any gradient updates—a phenomenon known as in-context learning (ICL; Brown et al., 2020). ICL, also referred to as few-shot learning, is a fundamental mechanism in LMs (Lampinen et al., 2025). Therefore, understanding the nature and generalization limits of ICL during the the next-token prediction pretraining process, is crucial. Additionally, understanding the nature of ICL is central to the safe and secure deployment of generative AI, as it has been argued that since increased scale alone is known to unlock new abilities in LMs, further scaling up has the potential to unlock latent hazardous abilities which carry important security and safety implications. Simultaneously, Lu et al. (2024) provide strong correlations between the ability of a model to solve tasks in few-shot ICL and their generalizability on novel tasks upon instruction tuning. Therefore, understanding the fundamental mechanisms of ICL has broad implications, including safety, security, and the capabilities of user-facing, instruction-tuned LMs.
Despite its importance, the current understanding of ICL in LMs remains limited and is often clouded by contradictory claims and findings, unlike our understanding of traditional in-weight learning via gradient descent. Specifically, researchers continue to debate whether ICL is functionally equivalent to gradient descent (Von Oswald et al. (2023) vs. Deutch et al. (2024)), and whether ICL is primarily regurgitation data and patterns encountered during pretrained or genuine generalization—even to the point of implementing symbolic algorithms (Golchin et al. (2024) vs. Elhage et al. (2021)), inter alia. A significant majority of prior attempts to investigate the characteristics of ICL focus on either Transformers trained on purely synthetic data (Edelman et al., 2024; Singh et al., 2023) or mechanism construction within toy-sized Transformers (Elhage et al., 2021; Olsson et al., 2022); as such, it is hard to scale these findings to LMs trained on real-world language data. Simultaneously, there has been a discrepancy in the tasks considered to characterize ICL; e.g., Golchin et al. (2024) consider different NLP benchmark tasks (with the possibility of leakage), while Olsson et al. (2022) define ICL more narrowly as an LM’s ability to replicate patterns from context. Intuitively, while the latter is a prerequisite for the former, several missing intermediary elements prevent a direct equivalence, leading to contradictory outcomes. This discrepancy within the scope of tasks to investigate language models’ ability (or inability) to infer abstract rules from the context and perform symbolic operations based on those rules has continued beyond few-shot task solving. Prior works confirm the existence of symbolic computation (Prakash et al., 2025; Yang et al., 2025) as well as its brittleness (Shojaee et al., 2025; Mirzadeh et al., 2024). However, the tasks these two groups consider are of completely different types, and subsequently, one cannot draw any generalized conclusion.
@@
We find that when provided with in-context examples, LMs trained on real-world corpora display a peculiar mix of generalized pattern-matching abilities on the one hand, and a dependence on pattern statistics on the other. Specifically, they can detect and follow patterns presented in-context using random token orderings—and therefore are unlikely to have memorized them during pre-training—refuting prior claims by Golchin et al.’s (2024). However, this pattern-learning ability declines when statistically rarer tokens are used to construct the patterns or when the pattern statistics in the examples are obfuscated, directly contradicting prior claims by Olsson et al.’s (2022), who describe ICL as symbolic algorithm execution. These two findings significantly refine our understanding of the mechanistic basis for ICL in LM, with implications to interpretability and circuit discovery (Wang et al., 2022; Conmy et al., 2023; Yu et al., 2024) in LMs, as contemporary approaches often operate under the assumption that symbolic functions are being implemented within the models’ internals.
@@
Upon investigating the model internals, we identify an aspect of the development that remains stable across different ICL characteristics. Specifically, we show that the dimensionality of the subspace of the residual stream (Elhage et al., 2021) used by the model to “carry” the answer token follows a consistent pattern: an initial rise in subspace allocation, corresponding to a “general” development of ICL competence, followed by a decay that saturates over training. Moreover, development of ICL competence through pretraining correlates with the model learning to use a common subspace across tasks. Both of these developments follow a predictable pattern, serving as a predictable reference variable for “occasionally abrupt” development of ICL competence. All of these findings bear importance to the notion of “emergent abilities of Large Language Models” introduced by Wei et al. (2022) and Ganguli et al. (2022) which are also related to ICL as described by Olsson et al. (2022); Lu et al. (2024): the predictability, or the lack thereof, of ICL has direct relevance to the possibility of other abilites being unexpectedly unlocked with further increase in scale, with significant implications to AI security.
@@ 2 Related Work
Brown et al. (2020) first empirically demonstrated that LMs, trained purely on next-token prediction (a.k.a. the language modeling task), can learn to perform a task by observing a few demonstrations in the input prompt. In this work, we coin the term ICL competence to refer to the general prowess of the LM in performing ICL, quantified by the model’s performance on various ICL tasks. Our notion of ICL competence is much similar to Lampinen et al. (2025) who argue that ICL encapsulate a broad range of tasks beyond supervised few-shot learning. The number of studies showcasing what tasks LMs can do from ICL examples is too many to discuss individually within the scope of this paper; instead, we refer to the elaborate survey by Dong et al. (2024a). Min et al. (2022) attempted to disentangle the factors influencing an LM’s ICL competence by perturbing the input-output space, subsequently claiming that LMs are often able to perform in-context classification correctly even when the labels are randomly permuted in the context. Wei et al. (2023) showcased that semantic priors of input-label mapping play a vital role in smaller LMs, while large LMs can perform ICL with semantically unrelated labels. Shi et al. (2023) explained this effect of scale on ICL competence by larger LMs’ ability to process more in-context features. These works remain primarily focused on the static abilities of the trained model, without investigating how these abilities might have arised via large-scale pretraining. In a somewhat orthogonal direction, Akyürek et al. (2024) investigated ICL competence through the lens of formal language modeling.
@@
A distinct line of attempts to understand ICL presents constructions of Transformer attention that can simulate gradient descent to argue that Transformers can implement certain classes of learning algorithms implicitly as ICL (Akyürek et al., 2022; Von Oswald et al., 2023). (Deutch et al., 2024) argued against this expressivity-based line of reasoning and pointed towards the distributional properties of training data that possibly dictate the development of ICL. This debate has been formalized to the question of whether ICL and gradient descent are equivalent or not (Dai et al., 2023; Mahdavi et al., 2024; Deutch et al., 2024). Prior work by Chan et al. (2022) has also pointed towards the data-distributional sources of ICL. Singh et al. (2023) demonstrated the transient nature of ICL in small Transformer models trained on synthetic data. (Chan et al., 2024) confirmed this transient nature and provided theoretical results that explain the learning dynamics of ICL. These works primarily identify the long-tailed distribution of the data as being the driver of ICL, whereas Bratulić et al. (2025) claimed that conceptual repetitions in the pretraining data are the primary determiner of ICL. Most of these prior investigations primarily rely on toy-sized Transformers and/or synthetic pretraining data. On the other hand, we focus on Transformers of a wide range of parameter scales, trained on real web-scale natural language data. To the best of our knowledge, Wang et al. (2024) comes as the only similar prior work in this direction; however, they sought to identify the dynamics of task recognition from ICL examples vs. task learning as they progress through pretraining.
@@
Elhage et al. (2021) and Olsson et al. (2022) were among the first to demonstrate that LMs can consistently complete previously seen patterns in-context: [A][B]...[A]→[B]. They conjectured that certain internal model mechanisms (specifically, induction heads) form to perform this pattern-completion task as a symbolic algorithm. This hypothesis has since underpinned mechanistic interpretability research, which assumes LMs develop internal mechanisms akin to symbolic algorithms and tasks researchers with identifying them. It has inspired projects like Tracr (Lindner et al., 2023), which compiles symbolic algorithms into transformer weights, and circuit discovery methods (Wang et al., 2022; Conmy et al., 2023) to uncover them. Cho et al. (2024) recently identified the mechanistic circuitry of ICL as a three-step process. Upon analysis on synthetically trained Transformers, Park et al. (2024) found that the development of ICL is a juxtaposition of multiple algorithms and their completion, and any claim of ‘singular’ ICL mechanism should be taken with caution. In a somewhat less fine-grained method of interpreting ICL, the notion of function vectors or task vectors (Todd et al., 2023) has been proposed. Yin & Steinhardt (2025) identified attention heads that construct function vectors and show that 1) these heads are formed from induction heads upon further pretraining, and 2) these heads are the fundamental driver of ICL-specific circuits and not induction heads, as presumed by Elhage et al. (2021). More recently, Yang et al. (2025) hypothesise a three‑stage symbolic processing circuit in LLMs for abstract rule‑induction reasoning: abstraction, pattern induction, and retrieval. Wu et al. (2025) also observe a similar process for variable binding. In line with the school of mechanistic interpretability, our work validates the algorithmic nature of ICL, but puts significant limitations on the notion as well. While we seek to identify signals internal to the model that implicate ICL and its development with training and parameter scaling (Section 6), we do not focus on recovering the exact neural algorithmic implementation of ICL.
@@
Table 1: An overview of our six proposed tasks: lsc, lscg, wc, wi, $t$, and cf. In the examples, the prompt text is highlighted in green, and the expected target token is highlighted in orange.
@@
Literal Sequence Copying (LSC): We start with providing a systematic empirical evaluation of the literal sequence copying phenomenon identified by Olsson et al. (2022). In particular, we aim to test whether the model can generate T when prompted with P T R P, where T is a target token and P and R are sequences of randomly sampled tokens of length ||P|| and ||R||, respectively.
@@
Task Configurations: Pattern length (||P||): the number of tokens in the literal sequence pattern prefix (P). Random gap length (||R||): the number of tokens in the random gap sequence (R).
@@
[…]
Task Configurations: Pattern length (||P||): the number of tokens in the literal sequence pattern prefix (P). Random gap length (||R||): the number of tokens in the random gap sequence (R). Random gap in pattern length (||G||): the number of tokens in the random gap sequence in the pattern with random gap (U or U).
@@
Task Configurations: Number of features (||F||): the number of tokens to identify in sequence. Number of labels (||L||): the number classification labels. Number of distractors (||D||): the number of distractor tokens in the classification sequence.
@@
Example: If we switch the capital of Canada and Germany, then Canada’s capital is Berlin and Germany’s capital is Ottawa